? Find a Frobenius series solution for x > 0. Use the small index. 4x2 y + 9xy + (-3x2 + 1) y = 0) 4x2 y + 7xy - (2x2 + 1

chose on of the two
Answer #1

Then htk 70 4&ry +9x41+ (-32 +1) 4 =0 since xeo is Ragulas singular par het &- & Antik be the solution -> 4x2 [124, (***) (kcase I k=1 An= 3 Anos (n-1) cant) tl A2 = 300 - 3 Ao lo 1x9+1 Az = 0 3 A2 3x 17+1 mie Ho = 9 520 Then ne Y, CA) = J I Hot A R
Answer #2

3. Find a Frobenius senies solution 30. use the small index. 404 tary +91 - 1) = 0 - 1 0 443-tang-(2x++)40 - Theorem a at lHonce the use small index I 40 (8-1) 4gr +17662er con + (7-1+9(ne) 11] Chung the lowest o the coerhonnt of Equating to zero* |(141-16 417 em - 36m2.30 - nu v-XX] cn - 36.220 - 46049 - 1) (-3 C. 2o. 2) [ 4n 13 cm - 3 (n.220 = c = 3 (n. 2 na (42137 2